# Expansion Span: Combining Fading Memory and Retrieval in Hybrid State Space Models

### v1

@@ Abstract
The ‘state’ of State Space Models (SSMs) represents their memory, which fades exponentially over an unbounded span. By contrast, Attention-based models have ‘eidetic’ (i.e., verbatim, or photographic) memory over a finite span (context size). Hybrid architectures combine State Space layers with Attention, but still cannot recall the distant past and can access only the most recent tokens eidetically. Unlike current methods of combining SSM and Attention layers, we allow the state to be allocated based on relevancy rather than recency. In this way, for every new set of query tokens, our models can ‘eidetically’ access tokens from beyond the Attention span of current Hybrid SSMs without requiring extra hardware resources. We describe a method to expand the memory span of the hybrid state by ‘reserving’ a fraction of the Attention context for tokens retrieved from arbitrarily distant in the past, thus expanding the eidetic memory span of the overall state. We call this reserved fraction of tokens the ‘expansion span,’ and the mechanism to retrieve and aggregate it ‘Span-Expanded Attention’ (SE-Attn). To adapt Hybrid models to using SE-Attn, we propose a novel fine-tuning method that extends LoRA to Hybrid models (HyLoRA) and allows efficient adaptation on long spans of tokens. We show that SE-Attn enables us to efficiently adapt pre-trained Hybrid models on sequences of tokens up to 8 times longer than the ones used for pre-training. We show that HyLoRA with SE-Attn is cheaper and more performant than alternatives like LongLoRA when applied to Hybrid models on natural language benchmarks with long-range dependencies, such as PG-19, RULER, and other common natural language downstream tasks.
@@ 1 Introduction
State Space Models are able to process sequences with an unbounded number of tokens by maintaining a fixed-size state. However, this state is lossy and information about early tokens ‘fades’ as more inputs are processed. In contrast, Transformer models have a state determined by the number of tokens in their input sequence and are able to access information from all past tokens in their context ‘eidetically.’ However, they do so at the cost of extra compute and memory. Recent Hybrid models [38, 12, 8] augment SSMs with Attention layers in an effort to counteract SSMs’ ‘fading’ memory. However, Attention layers only aggregate information by recency (i.e., they process the keys and values of the most recent tokens up to hardware limitations). Hence, any token that lies beyond the Attention’s limited span can only be approximately recalled through the SSMs’ state. This limits the effectiveness of Hybrid SSMs when applied to long sequences of tokens, especially when compute and memory resources are limited. In fact, while recent pre-trained Hybrid models have demonstrated strong performance on common downstream and recall-intensive language tasks, they are trained following Transformer’s pre-training practice: they are either trained on a fixed, relatively short window of tokens when compute is limited—which often is set to 2048 as in LLaMA1 [31], or 4096 as in LLaMA2 [32]—or are progressively fine-tuned on longer sequences after the pre-training stage [22, 12, 9].
@@
In this work, we modify Hybrid SSMs’ Attention layers to allow their state to be allocated by relevancy rather than recency, allowing our models to retrieve information that would otherwise fall outside their Attention span. To do so, we propose Span-Expanded Attention (SE-Attn), a selection mechanism that complements the ‘fading’ memory of SSMs with ‘eidetically’ retrieved tokens.
@@
Our method allows long context fine-tuning of pre-trained Hybrid models on sequences longer than the ones used for pre-training without significant compute/memory overhead. While Transformer-based fine-tuning strategies for processing long sequences can also be applied to Hybrid models, they typically work under the assumption that information in the Attention layers is aggregated by recency and are thus more expensive as sequences get longer. For example, extending a LLaMA model from a 2k to 8k context size can require up to 32 A100 GPUs [2]. While there exist efficient LoRA-based methods [17] to bring this cost down, we show that Transformer-based methods tend to not work as well for Hybrid models as they do for Transformers. We show that this is mostly due to the fact that combining LoRA (on Attention layers) with SSM layers is not expressive enough to model long-range dependencies beyond the pre-training context size.
@@
To overcome such limitations, we propose to modify the standard LoRA recipe to fit the Hybrid model structure better. Namely, we employ a fine-tuning strategy similar to LoRA+ [3] on Attention layers, while allowing the SSMs’ recurrent parameters to be adapted as well. In particular, building on previous observations [38, 37], we also adapt the 1D convolutional layers, which we empirically validate provide the best results at a reduced computational cost.
We make the following contributions:
  1. 1.
    We propose SE-Attn, a novel selection mechanism that ‘reserves’ a fraction of the context of Hybrid SSMs’ Attention layers for tokens that are retrieved from the past based on their relevance to the current query.
  2. 2.
    We empirically show that Transformer-based methods for long-context fine-tuning are sub-optimal when applied to Hybrid SSMs. Therefore, we propose HyLoRA, an extension of LoRA+ [3] for efficient and effective fine-tuning of Hybrid SSMs.
  3. 3.
    We show that combining SE-Attn and HyLoRA on Hybrid models reliably extends their context size up to 8×8$\times$8 × their pre-training size. Further, we show improved performance over existing adaptation methods on common natural language tasks in the LM Harness Evaluation suite [11] and long context tasks in RULER [16].
@@ 2 Background and Related Work
Attention on long contexts. Since the introduction of Self-Attention [33], a significant amount of research has been made to reduce its quadratic cost in the sequence length at training time. To achieve this, state of the art methods usually approximate the Attention mechanism with sparse/linearized versions. For example, Reformer [21] uses locality-sensitive-hashing to group tokens with similar embeddings, allowing the model to only attend to a subset of tokens rather than the entire sequence. Longformer [1] combines local Attention, whereby each token is restricted to attending only to a window of neighboring tokens, and global Attention, whereby a few tokens can attend to all tokens. Meanwhile, Linformer [35] uses projection matrices to project keys and values into a lower-dimensional space along the sequence dimension.
@@
Different from these approximate Attention methods, other works have proposed to endow Transformers models with ‘compressed’ memory tokens that are updated dynamically and causally over sliding windows on entire sequence chunks. For example, Transformer-XL [4] and Infini-Attention [24] segment an input sequence into chunks and process them sequentially while maintaining a complementary set of tokens whose purpose is to summarize the older ones. In contrast to these works, our Attention retrieves relevant tokens ‘eidetically,’ i.e., we retrieve tokens rather than maintain a compressed representation of the past. Most similar to our method is Landmark Attention [23], which inserts landmark tokens into the input at fixed block intervals and trains these tokens to act as summaries of their corresponding blocks via a grouped softmax attention; these summary tokens are then used to index and retrieve relevant input tokens when processing future segments. Our Attention aims to integrate retrieval natively, without the need for external landmark tokens or complex Softmax procedures.
@@
State Space Models. While a great effort has been made to improve the efficiency of Transformer models, a recent line of work has explored efficient alternative ‘linear’ architectures. In particular, State Space Models (SSMs) [13, 15, 14, 36, 30] have emerged as promising competitors to Transformer models due to their efficient scaling and strong empirical performance. Numerous variations of SSMs have been proposed, some closely resembling Linear Attention [30] or Linear-Time Invariant dynamical systems [15], while others introduce novel adaptive/gated state updates [36, 13, 6, 25]. Despite their differences, all follow the same basic working principle that is inspired by classical state space models [19]: they process the input sequence by maintaining a fixed-size state which acts as a compressed (lossy) representation of all the processed tokens. However, when implemented in hardware, the state must have finite precision and therefore ‘fades’ as more samples are processed. To overcome this limitation, the most successful SSMs are typically hardware-aware implementations that efficiently utilize modern GPUs/TPUs. These implementations use highly parallelizable and scalable primitives, such as associative scans [13, 8], chunking mechanisms [6, 36], and techniques that avoid the materialization of the entire state in slow high bandwidth memory [13].
@@
Hybrid State Space Models. While extending the recurrent state in SSMs layers has led to performant models, they typically fall short on tasks that require recalling information from the distant past [34, 18]. Hybrid State Space Models have been introduced to solve this limitation and are typically designed to complement the SSMs’ ‘fading’ state with Attention layers [7, 8, 22, 12]. While early architectures simply stack SSMs and Attention layers with different blending ratios [34, 13, 6] or replace full Attention layers with Sliding Window Attention [8], more performant and sophisticated hybrid architectures have been recently proposed [12, 38]. In particular, B’MOJO [38] complements SSMs’ fading state with a form of ‘eidetic’ memory whereby SSM layers are mixed with Sliding Window Attention and tokens in the window can attend to a set of tokens from the past that are deemed difficult to predict using an asynchronous causal selection mechanism. Similarly, in our work, we mix SSM layers with an Attention layer that operates over chunks of tokens, but we enable these chunks to dynamically retrieve tokens from the past that are most relevant to the chunk’s tokens.
@@ 3 Span-Expanded Attention
Figure 1: Span-Expanded Attention (SE-Attn) overview. SE-Attn is a sparse Attention mechanism used to expand the memory span of Hybrid SSMs. We do so by reserving a fraction of the Attention context for tokens retrieved arbitrarily far back in the past, and use summary tokens to efficiently look back with a reduced compute cost. We call this reserve the ‘expansion span,’ and we populate it with blocks of previous tokens (memory blocks). When new tokens arrive, a similarity-based search compares the queries with past memory blocks to decide which memory blocks are most relevant. Then, these retrieved memory blocks are jointly processed with the queries via Attention. While the final Attention mechanism always processes a fixed number of tokens, it can have a longer span since tokens from arbitrarily far back in the past can be retrieved.
Our goal is to enable pre-trained Hybrid SSMs to accurately process sequences with a larger number of tokens than were used for pre-training. For a sequence of $L$ tokens with model dimension $d_{\textrm{model}}$, the computational complexity of standard Self-Attention is $O(d_{\textrm{model}}L^{2})$. For very large $L$, computing Self-Attention can be prohibitively expensive. Training models with a large $L$ is particularly challenging, as training has the additional cost of computing gradients, further limiting memory resources. To address this, we propose Span-Expanded Attention (SE-Attn), a drop-in replacement for standard Self-Attention in Hybrid SSMs. To train SE-Attn, we propose HyLoRA, a variant of LoRA+ [3] specifically tailored to Hybrid models. A schematic of our SE-Attn is provided in Figure 1.

### v2

@@ Abstract
The “state” of State Space Models (SSMs) represents their memory, which fades exponentially over an unbounded span. By contrast, Attention-based models have “eidetic” (i.e., verbatim, or photographic) memory over a finite span (context size). Hybrid architectures combine State Space layers with Attention, but still cannot recall the distant past and can access only the most recent tokens eidetically. Unlike current methods of combining SSM and Attention layers, we allow the state to be allocated based on relevancy rather than recency. In this way, for every new set of query tokens, our models can “eidetically” access tokens from beyond the Attention span of current Hybrid SSMs without requiring extra hardware resources. We introduce a method to expand the memory span of the hybrid state by “reserving” a fraction of the Attention context for tokens retrieved from arbitrarily distant in the past, thus expanding the eidetic memory span of the overall state. We call this reserved fraction of tokens the “expansion span,” and the mechanism to retrieve and aggregate it “Span-Expanded Attention” (SE-Attn). To adapt Hybrid models to using SE-Attn, we propose a novel fine-tuning method that extends LoRA to Hybrid models (HyLoRA) and allows efficient adaptation on long spans of tokens. We show that SE-Attn enables us to efficiently adapt pre-trained Hybrid models on sequences of tokens up to 8 times longer than the ones used for pre-training. We show that HyLoRA with SE-Attn is cheaper and more performant than alternatives like LongLoRA when applied to Hybrid models on natural language benchmarks with long-range dependencies, such as PG-19, RULER, and other common natural language downstream tasks.
keywords:
Hybrid State Space Models, SSMs, Attention, Sparse Attention, LoRA.
@@ 1 Introduction
State Space Models are able to process sequences with an unbounded number of tokens by maintaining a fixed-size state. However, this state is lossy and information about early tokens “fades” as more inputs are processed. In contrast, Transformer models have a state determined by the number of tokens in their input sequence and are able to access information from all past tokens in their context “eidetically.” However, they do so at the cost of extra compute and memory.
@@
Recent Hybrid models [Zancato et al.(2024)Zancato, Seshadri, Dukler, Golatkar, Shen, Bowman, Trager, Achille, and Soatto, Glorioso et al.(2024b)Glorioso, Anthony, Tokpanov, Whittington, Pilault, Ibrahim, and Millidge, De et al.(2024)De, Smith, Fernando, Botev, Cristian-Muraru, Gu, Haroun, Berrada, Chen, Srinivasan, et al.] augment SSMs with Attention layers in an effort to counteract SSMs’ “fading” memory. However, Attention layers only aggregate information by recency (i.e., they process the keys and values of the most recent tokens up to hardware limitations). Hence, any token that lies beyond the Attention’s limited span can only be approximately recalled through the SSMs’ state. This limits the effectiveness of Hybrid SSMs when applied to long sequences of tokens, especially when compute and memory resources are limited. In fact, while recent pre-trained Hybrid models have demonstrated strong performance on common downstream and recall-intensive language tasks, they are trained following Transformers’ pre-training practice: they are either trained on a fixed, relatively short window of tokens when compute is limited—which is often set to 2048 as in LLaMA1 [Touvron et al.(2023a)Touvron, Lavril, Izacard, Martinet, Lachaux, Lacroix, Rozière, Goyal, Hambro, Azhar, et al.], or 4096 as in LLaMA2 [Touvron et al.(2023b)Touvron, Martin, Stone, Albert, Almahairi, Babaei, Bashlykov, Batra, Bhargava, Bhosale, et al.]—or are progressively fine-tuned on longer sequences after the pre-training stage [Lieber et al.(2024)Lieber, Lenz, Bata, Cohen, Osin, Dalmedigos, Safahi, Meirom, Belinkov, Shalev-Shwartz, et al., Glorioso et al.(2024b)Glorioso, Anthony, Tokpanov, Whittington, Pilault, Ibrahim, and Millidge, Dubey et al.(2024)Dubey, Jauhri, Pandey, Kadian, Al-Dahle, Letman, Mathur, Schelten, Yang, Fan, et al.].
@@
In this work, we modify Hybrid SSMs’ Attention layers to allow their state to be allocated by relevancy rather than recency, allowing our models to retrieve information that would otherwise fall outside their Attention span. To do so, we propose Span-Expanded Attention (SE-Attn), a selection mechanism that enables eidetic retrieval of tokens from the past. While SE-Attn can be applied to both Hybrid SSMs and Transformers, we focus on hybrid architectures where this retrieval capability naturally complements the fading memory of SSMs: while SSMs efficiently process long sequences with their state-based memory, SE-Attn selectively preserves important information that would otherwise fade in the SSM state.
@@
Our method allows long context fine-tuning of any pre-trained Hybrid model, such as Mamba-2-Hybrid [Dao and Gu(2024)], Jamba [Lieber et al.(2024)Lieber, Lenz, Bata, Cohen, Osin, Dalmedigos, Safahi, Meirom, Belinkov, Shalev-Shwartz, et al.], and Zamba [Glorioso et al.(2024a)Glorioso, Anthony, Tokpanov, Golubeva, Shyam, Whittington, Pilault, and Millidge], on sequences longer than the one used for pre-training without significant compute/memory overhead. While Transformer-based fine-tuning strategies for processing long sequences can also be applied to Hybrid models, they typically assume that information in the Attention layers is aggregated by recency and are thus more expensive as sequences get longer. For example, extending a LLaMA model from a 2k to 8k context size can require up to 32 A100 GPUs [Chen et al.(2023)Chen, Wong, Chen, and Tian]. While there exist efficient LoRA-based methods [Hu et al.(2022)Hu, yelong shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen] that lower this cost, we show that Transformer-based methods tend to not work as well for Hybrid models as they do for Transformers. We show that this is mostly due to the fact that combining LoRA (on Attention layers) with SSM layers is not expressive enough to model long-range dependencies beyond the pre-training context size.
@@
To overcome such limitations, we propose to modify the standard LoRA recipe to fit the Hybrid model structure better. Namely, we employ a fine-tuning strategy similar to LoRA+ [Chen et al.(2024)Chen, Qian, Tang, Lai, Liu, Han, and Jia] on Attention layers, while allowing the SSMs’ recurrent parameters to be adapted as well. In particular, building upon previous observations [Zancato et al.(2024)Zancato, Seshadri, Dukler, Golatkar, Shen, Bowman, Trager, Achille, and Soatto, Yang et al.(2024b)Yang, Wang, Zhang, Shen, and Kim], we also adapt the 1D convolutional layers; we empirically demonstrate that this adaptation yields the best results at a reduced cost.
We make the following contributions: (1) We propose SE-Attn, a novel selection mechanism that “reserves” a fraction of the context of Hybrid SSMs’ Attention layers for tokens that are retrieved from the past based on their relevance to the current query. (2) We empirically show that Transformer-based methods for long-context fine-tuning are sub-optimal when applied to Hybrid SSMs. Therefore, we propose HyLoRA, an extension of LoRA+ [Chen et al.(2024)Chen, Qian, Tang, Lai, Liu, Han, and Jia] for efficient and effective fine-tuning of Hybrid SSMs. (3) We show that combining SE-Attn and HyLoRA on Hybrid models reliably extends their context size up to 8×8$\times$8 × their pre-training size. Further, we show improved performance over existing adaptation methods on common natural language tasks in the LM Harness Evaluation suite [Gao et al.(2024)Gao, Tow, Abbasi, Biderman, Black, DiPofi, Foster, Golding, Hsu, Le Noac’h, Li, McDonell, Muennighoff, Ociepa, Phang, Reynolds, Schoelkopf, Skowron, Sutawika, Tang, Thite, Wang, Wang, and Zou] and long context tasks in RULER [Hsieh et al.(2024)Hsieh, Sun, Kriman, Acharya, Rekesh, Jia, Zhang, and Ginsburg].
@@ 2 Background and Related Work
Attention on long contexts. Since the introduction of Self-Attention [Vaswani et al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin], a significant amount of research has been made to reduce its quadratic cost in the sequence length at training time. To achieve this, state of the art methods usually approximate the Attention mechanism with sparse/linearized versions. For example, Reformer [Kitaev et al.(2020)Kitaev, Kaiser, and Levskaya] uses locality-sensitive-hashing to group tokens with similar embeddings, allowing the model to only attend to a subset of tokens rather than the entire sequence. Other works have proposed to endow Transformers models with “compressed” memory tokens that are updated dynamically and causally over sliding windows on entire sequence chunks. For example, Transformer-XL [Dai(2019)] and Infini-Attention [Munkhdalai et al.(2024)Munkhdalai, Faruqui, and Gopal] segment an input sequence into chunks and process them sequentially while maintaining a complementary set of tokens whose purpose is to summarize the older ones. In contrast to these works, our SE-Attn retrieves relevant tokens “eidetically,” i.e., we retrieve tokens rather than maintain a compressed representation of the past. Most similar to our method is Landmark Attention [Mohtashami and Jaggi(2023)], which inserts landmark tokens into the input at fixed block intervals and trains these tokens to act as summaries of their corresponding blocks via a grouped softmax attention; these summary tokens are then used to index and retrieve relevant input tokens when processing future segments. Our SE-Attn aims to integrate retrieval natively, without the need for external landmark tokens or complex Softmax procedures.
@@
State Space Models. While a great effort has been made to improve the efficiency of Transformer models, a recent line of work has explored efficient alternative “linear” architectures. In particular, State Space Models (SSMs) [Gu and Dao(2023), Gu et al.(2021)Gu, Johnson, Goel, Saab, Dao, Rudra, and Ré, Gu et al.(2022)Gu, Goel, and Re, Yang et al.(2024a)Yang, Wang, Shen, Panda, and Kim, Sun et al.(2023)Sun, Dong, Huang, Ma, Xia, Xue, Wang, and Wei] have emerged as promising competitors to Transformer models due to their efficient scaling and strong empirical performance. Numerous variations of SSMs have been proposed, some closely resembling Linear Attention [Sun et al.(2023)Sun, Dong, Huang, Ma, Xia, Xue, Wang, and Wei] or Linear-Time Invariant dynamical systems [Gu et al.(2021)Gu, Johnson, Goel, Saab, Dao, Rudra, and Ré], while others introduce novel adaptive/gated state updates [Yang et al.(2024a)Yang, Wang, Shen, Panda, and Kim, Gu and Dao(2023), Dao and Gu(2024), Orvieto et al.(2023)Orvieto, Smith, Gu, Fernando, Gulcehre, Pascanu, and De]. Despite their differences, all follow the same basic working principle that is inspired by classical state space models [Kalman(1960)]: they process the input sequence by maintaining a fixed-size state which acts as a compressed (lossy) representation of all the processed tokens. When implemented in hardware, the state must have finite precision and therefore “fades” as more samples are processed. To overcome this limitation, the most successful SSMs are typically hardware-aware implementations that efficiently utilize modern GPUs/TPUs. These implementations use highly parallelizable and scalable primitives, such as associative scans [Gu and Dao(2023), De et al.(2024)De, Smith, Fernando, Botev, Cristian-Muraru, Gu, Haroun, Berrada, Chen, Srinivasan, et al.], chunking mechanisms [Dao and Gu(2024), Yang et al.(2024a)Yang, Wang, Shen, Panda, and Kim], and techniques that avoid the materialization of the entire state in slow high bandwidth memory [Gu and Dao(2023)].
Figure 1: Span-Expanded Attention (SE-Attn) overview. SE-Attn is a Sparse Attention mechanism used to expand the memory span of Hybrid SSMs. Left: SE-Attn works by reserving a fraction of the Attention context for tokens retrieved arbitrarily far back in the past. We call this reserve the “expansion span,” and we populate it with blocks of previous tokens (“memory blocks”). When new tokens arrive, a similarity-based search compares the queries with past memory blocks—represented as summary tokens—to retrieve relevant memory blocks. Then, these retrieved memory blocks are jointly processed with the queries via Attention. While the final Attention mechanism always processes a fixed number of tokens, it can have a longer span since it retrieves tokens from arbitrarily far back in the past. Right: Retrieving tokens from the past yields a sparse Attention pattern.
@@
Hybrid State Space Models. While extending the recurrent state in SSMs layers has led to performant models, they typically fall short on tasks that require recalling information from the distant past [Waleffe et al.(2024)Waleffe, Byeon, Riach, Norick, Korthikanti, Dao, Gu, Hatamizadeh, Singh, Narayanan, et al., Jelassi et al.(2024)Jelassi, Brandfonbrener, Kakade, and Malach]. Hybrid State Space Models have been introduced to solve this limitation and are typically designed to complement the SSMs’ “fading” state with Attention layers [Dao and Gu(2024), De et al.(2024)De, Smith, Fernando, Botev, Cristian-Muraru, Gu, Haroun, Berrada, Chen, Srinivasan, et al., Lieber et al.(2024)Lieber, Lenz, Bata, Cohen, Osin, Dalmedigos, Safahi, Meirom, Belinkov, Shalev-Shwartz, et al., Glorioso et al.(2024b)Glorioso, Anthony, Tokpanov, Whittington, Pilault, Ibrahim, and Millidge]. While early architectures simply stack SSMs and Attention layers with different blending ratios [Waleffe et al.(2024)Waleffe, Byeon, Riach, Norick, Korthikanti, Dao, Gu, Hatamizadeh, Singh, Narayanan, et al., Gu and Dao(2023), Dao and Gu(2024)] or replace full Attention layers with Sliding Window Attention [De et al.(2024)De, Smith, Fernando, Botev, Cristian-Muraru, Gu, Haroun, Berrada, Chen, Srinivasan, et al.], more performant and sophisticated hybrid architectures have been recently proposed [Glorioso et al.(2024b)Glorioso, Anthony, Tokpanov, Whittington, Pilault, Ibrahim, and Millidge, Zancato et al.(2024)Zancato, Seshadri, Dukler, Golatkar, Shen, Bowman, Trager, Achille, and Soatto]. In particular, B’MOJO [Zancato et al.(2024)Zancato, Seshadri, Dukler, Golatkar, Shen, Bowman, Trager, Achille, and Soatto] complements SSMs’ fading state with a form of “eidetic” memory whereby SSM layers are mixed with Sliding Window Attention and tokens in the window can attend to a set of tokens from the past that are deemed difficult to predict using an asynchronous causal selection mechanism. Similarly, in our work, we mix SSM layers with an Attention layer that operates over chunks of tokens, but we enable these chunks to dynamically retrieve tokens from the past that are most relevant to the chunk’s tokens, rather than tokens selected a priori as in B’MOJO.
@@ 3 Span-Expanded Attention
Our goal is to enable pre-trained Hybrid SSMs to accurately process sequences longer than the ones seen during pre-training. For a sequence of $L$ tokens with model dimension $d_{\textrm{model}}$, the computational complexity of standard Self-Attention is $O(d_{\textrm{model}}L^{2})$. For very large $L$, computing Self-Attention can be prohibitively expensive. Training models with a large $L$ is particularly challenging, as training has the additional cost of computing gradients, further limiting memory resources. To address this, we propose Span-Expanded Attention (SE-Attn), a drop-in replacement for standard Self-Attention in Hybrid SSMs. To train SE-Attn, we propose HyLoRA, a variant of LoRA+ [Chen et al.(2024)Chen, Qian, Tang, Lai, Liu, Han, and Jia] specifically tailored to Hybrid models. A schematic of our SE-Attn is provided in Figure 1.